OpenStudy (anonymous):

Find the sum 1/2 + 1/6=

OpenStudy (anonymous):

I Keep getting like 0.666666666666666666666666667

OpenStudy (anonymous):

3 + 1/6 = 4/6 = 2/3 in terms of fractions.

OpenStudy (anonymous):

The answer that you got is right.